OpenStudy (anonymous):

PLEASE HELP!! factor the following completely: 8x^4-6x^2-27

OpenStudy (anonymous):

Take \(x^2 = t\) and then factor.

OpenStudy (anonymous):

\[8t^2 - 6t -27=0\]

OpenStudy (chaise):

x^2 = t 8t^2 - 6t-27 (2t+3)(4t-9) Replace your t with x^2 (2x^2+3)(4x^2-9)

OpenStudy (anonymous):

can you please show me and explain the steps please? here is the equation again: \[8x^{4}-6x ^{2}-27\]

OpenStudy (chaise):

I will explain it for you :) 8x^4-6x^2-27 If we let x^2 = t we are able to factorise this equation VERY easily. t = x^2 t^2 = x^4 8t^2 - 6t - 27 = 6x^4 - 6x^2 - 27; according to these restrictions> 8t^2 - 6t - 27 can be factorised to: (2t+3)(4t-9). Do you need help factorising? Now we have the factorised version, we can substitute x^2 instead of t in this equation we get: (2t+3)(4t-9) = (2x^2+3)(4x^2-9) Do you want me to write this into equation editor so it's a little clearer?

OpenStudy (anonymous):

thank you for trying to answer my question, yes, the equation editor would really help. andi need a little help with factorising.

OpenStudy (chaise):

\[8x^4 - 6x^2 - 27\] \[Let,t = x^2 and t^2 = x^4\] \[8t^2 - 6t - 27 = 6x^4 - 6x^2 - 27\] \[8t^2 - 6t - 27 = (2t+3)(4t-9)\] If we were to expand \[(2t+3)(4t-9)\] we would get: \[8t^2 - 6t - 27\] Factorising is opposite to expanding. So we are just going backward to simplify the question. \[(2t+3)(4t-9)\] Now we have this equation we can simply substitute back in our t = x^2 and t^2 = x^4 into the equation. \[(2x^2+3)(4x^2-9)\] This is the fully factorised version of your function. If you have trouble with factorising, have a look at some videos on factorising bionomials on youtube.

OpenStudy (anonymous):

thank you for your time and help. :D

OpenStudy (chaise):

Sorry if I made no sense at all.

OpenStudy (anonymous):

no, you made perfect sense :). i understand it now. ty.